OpenStudy (anonymous):

*Medal will be given* Change this radical to an algebraic expression with fractional exponents. 5√x^3

OpenStudy (gorv):

5*x^3/2=5x^1.5

OpenStudy (anonymous):

\[5\sqrt{x^3}=5{x^3}^{\times \frac{1}{2}}=5{x}^{\frac{3}{2}}\] @DjNck